### v1

# ValuesRAG: Enhancing Cultural Alignment Through Retrieval-Augmented Contextual Learning

Abstract
Cultural values alignment in Large Language Models (LLMs) is a critical challenge due to their tendency to embed Western-centric biases from training data, leading to misrepresentations and fairness issues in cross-cultural contexts. Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively. To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with in-context learning to integrate cultural and demographic knowledge dynamically during text generation. Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual. Subsequently, we curated several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries. ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided, highlighting ValuesRAG’s potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
1 Introduction
The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  Singh et al. (2024); Kharchenko et al. (2024); Hu et al. (2024). Predominantly trained on Western data sources Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts. These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024); Huang et al. (2024). Despite ongoing efforts to address these issues, existing strategies often fall short. While some countries have developed localized LLMs, such as China’s ERNIE Sun et al. (2021), South Korea’s HyperCLOVA Yoo et al. (2024), and the multilingual ChatGLM GLM et al. (2024), these models also exhibit biases inherited from their respective training datasets. As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024).
Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and the proposed ValuesRAG framework.
Recent studies have proposed several approaches, such as role-assignment approaches and few-shot learning techniques  Li et al. (2024a); AlKhamissi et al. (2024), to mitigate these cultural biases. However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and in-context learning to dynamically incorporate cultural knowledge during text generation (see Figure 1). Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al. (2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies. Specifically, we first generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel. After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected. Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies step-by-step reasoning grounded in the retrieved values to generate the final answer to the question.
We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al. (2024), (3) few-shot learning Choenni and Shutova (2024), and (4) a hybrid method combining (1) and (2). To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs. Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines, also in the ablation study where only summaries of values were provided without any additional information. Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications. Our study demonstrates ValuesRAG’s robust capabilities on a global scale, suggesting its applicability in aligning the values of diverse groups within a single country. ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al. (2024b). This, in turn, facilitates the creation of fairer and more effective policies. Moreover, NGOs can leverage ValuesRAG to develop LLMs that reflect specific value orientations while maintaining strong alignment with users’ values, thereby increasing their persuasive impact. This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
2 Related Work
2.1 Evaluation of LLMs’ Cultural Bias
Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue Tao et al. (2024). Unlike the more obvious safety concerns and social discrimination Liu et al. (2024) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data. Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al. (2024); Durmus et al. (2024). This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  Manvi et al. (2024). Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models Gallegos et al. (2024). For example, Webster et al.  Webster et al. (2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models, while Caliskan et al. Caliskan et al. (2017) pioneered the use of word embeddings as quantitative measures of bias. More recently, Karinshak et al. Karinshak et al. (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol. These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
Set Dataset Abbreviation Region Year Na VQb
Train Set World Values Survey WVS Global 2017–2022 97.2k 259
Test Sets European Values Study EVS Europe 2017 59.4k 211
The General Social Survey GSS North America 2021–2022 8.2k 44
Chinese General Social Survey CGSS East Asia 2021 8.1k 58
India Survey Dataset ISD South Asia 2019–2020 30.0k 33
AmericasBarometer LAPOP Latin America 2021 59.1k 48
Afrobarometer Afrobarometer Africa 2022 48.1k 144
Table 1: Overview of the datasets utilized in our study. The Train Set (WVS) includes global data collected between 2017 and 2022, providing the basis for generating cultural summaries of values and validation for our method. The Test Set consists of six region-specific surveys, each capturing socio-cultural information from distinct geographic areas and time frames. aN represents sample size in thousands (k). bVQ represents the number of values-related questions. cDQ represents the number of demographic features.
2.2 Mitigation of LLMs’ Cultural Bias
Methods like RLHF Shen et al. (2023); Ji et al. (2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al. (2024), as there are hundreds of countries and cultures in the earth. Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs. For instance, Tao et al. Tao et al. (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al. Masoud et al. (2024) developed a soft prompt tuning approach to mitigate bias. Also, Choenni and Shutova Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts. However, these approaches face significant limitations in fully capturing the complexity of cultural alignment. Tao et al.’s technique mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations. Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity. We, therefore, use these methods as baselines to benchmark our proposed approach.
3 Datasets
In this section, we first introduce the World Values Survey (WVS) as our training dataset (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies. Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
3.1 Train Dataset
WVS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al. (2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries. WVS is selected as our train dataset especially due to its numerous advantages:
1. Broad Recognition and Inclusiveness:
WVS is widely recognized and used by governments, social scientists, and international organizations in values studies. It covers 120 countries, representing 94.5% of the global population, ensuring broad geographic and cultural representation.
2. Expert-Designed and Accessible:
The dataset is designed by leading domain experts to conduct comprehensive surveys of values, ensuring reliability and relevance. It is publicly accessible, enabling reproducibility and transparency in research.
3. Effective Structure and Large Scale:
WVS has well-organized and comprehensive demographic questions, making it effective for retrieval tasks. Its large sample size (97,221 respondents) is also suitable for RAG tasks.
Since values evolve gradually over time, WVS is conducted in waves, with each wave occurring every five years. For our study, we utilized the most recent wave, spanning from 2017 to 2022.
The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions. The values questions span 13 topics, such as social trust, post-materialism, and political interest. We randomly selected 20% (52 questions) per topic for validation and used the remaining 80% (207 questions) for summary generation. The 31 demographic features, including country, sex, age, education, social class, and employment status, were used to generate demographic summaries for retrieval tasks.
3.2 Test Datasets
We selected six regional surveys to serve as test datasets based on the following criteria:
1. Demographic and Values Coverage:
The datasets provide demographic features closely aligned with WVS’s questions, along with sufficient values-related features to enable meaningful comparisons and analyses.
2. Temporal Proximity:
The datasets exhibit temporal proximity to WVS Wave 7 (2017–2022), allowing aligned comparisons and ensuring consistency across evaluations.
The regions in our test datasets were meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring the data accurately reflects the majority of the global population. All of them are publicly accessible, and are statistically representative at national or regional levels, which guarantees their reliability and validity.
To be specific, we selected European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region. For the United States, we selected the General Social Survey Davern et al. (2024), which is the most comprehensive social survey in the country. The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor. For India, where national survey data were largely inaccessible during our study, we used Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population. The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, was selected to represent Latin American countries, as it covers 32 countries across the region. Finally, Afrobarometer Afrobarometer (2023) was chosen as the representative dataset for Africa. A detailed summary of these datasets is presented in Table 1.
Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment, consisting of four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
4 Methodology
In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach. ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3). An overview of the ValuesRAG framework is provided in Figure 2.
4.1 Values and Demographic Summary Generation
To systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages. First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1). In parallel, topic-based summaries and demographic summaries are generated separately. For topic-based summaries, values-related QA sets are used to produce summaries for each topic, while demographic summaries are generated using demographic-related QA sets:
$T_{i}^{j}$ $=f_{\text{gen}}(\text{QA}_{i}^{j,1},\text{QA}_{i}^{j,2},\dots,\text{QA}_{i}^{j,N_{j}}),$ (1)
$D_{i}$ $=f_{\text{gen}}(\text{QA}_{i}^{\text{demo},1},\text{QA}_{i}^{\text{demo},2},\dots,\text{QA}_{i}^{\text{demo},K}).$
where $f_{\text{gen}}$ denotes the generative model, $T_{i}^{j}$ is the summary for topic $j$ of individual $i$, based on $N_{j}$ values-related QA pairs, and $D_{i}$ represents the demographic summary derived from $K$ demographic-related QA pairs. Finally, individual summaries are constructed by combining all topic summaries:
$\small S_{i}=f_{\text{gen}}(T_{i}^{1},T_{i}^{2},\dots,T_{i}^{M}),$ (2)
with $M$ denoting the total number of topics. The result, denoted as $S_{i}$, forms a comprehensive values summary for individual $i$. These generate summaries serve as structured references for retrieval in later stages and are also used to augment the validation set for evaluation.
4.2 Values Augmented Generation
Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating the validation question-answer data. For each validation question, we concatenate the corresponding individual’s values summary with the question itself, forming a context-rich input for the LLM:
$C_{i}=\text{concat}\left(S_{i},\text{Q}_{i}^{\text{val},k}\right)$ (3)
where $C_{i}$ represents the combined context, $S_{i}$ is the values summary for individual $i$, and $\text{Q}_{i}^{\text{val},k}$ is the $k$-th validation question. Next, we concatenate $C_{i}$ with the demographic summary $D_{i}$ to further enhance the context, enabling the generation of responses based on both values and demographic information:
$A_{i}=f_{\text{gen}}\left(C_{i},D_{i}\right)$ (4)
here, $A_{i}$ represents the answer generated by the function $f$, and $(C_{i},D_{i})$ embeds the augmented context $C_{i}$ and demographic information $D_{i}$ into a structured input format. Additionally, we utilize chain-of-thought prompting to enhance reasoning and emulate the behavior of the corresponding individual, ensuring responses that are contextually aligned with the values captured in the summaries and demographic characteristics.
4.3 Retrieval-based Values Alignment
To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval. The demographic data from both the train and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model. We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
Methods EVS GSS CGSS ISD LAPOP Africa Avg.
Baseline
Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
Role-Assignment + Few-Shot Learning 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
Ours
ValuesRAG($k=1$) 0.5960 0.7722* 0.5347* 0.6853* 0.4682 0.5904* 0.6078*
ValuesRAG($k=3$) 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
ValuesRAG($k=5$) 0.6051* 0.7706* 0.5301* 0.7016* 0.5061* 0.5905* 0.6173*
ValuesRAG($k=10$) 0.6020* 0.7380 0.5317* 0.7014* 0.4686 0.5680* 0.6016*
Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets, including the European Values Study (EVS), General Social Survey (GSS), Chinese General Social Survey (CGSS), Indian Survey Data (ISD), LAPOP, Afrobarometer, and the AmericasBarometer. $k$ indicates the number of summaries to be retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance. * denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p$ <0.05) over all baseline model(s).
$\small\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score. The top-100 embeddings with the highest similarity scores are initially selected as candidates. Subsequently, we apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates. The reranking process evaluates the semantic relevance of the initial candidates based on a scoring function:
$\small R^{\prime}_{k}=f_{\text{rerank}}\left(E_{\text{test}},E_{C_{j}}\right),\;C_{j}\in\{C_{1},C_{2},\dots,C_{100}\},\;k\leq 100$ (6)
where $R^{\prime}_{k}$ represents the reranked summary for the $k$-th candidate selected from the final top-$k$ results, $f_{\text{rerank}}$ is the reranking function, and $E_{C_{j}}$ denotes the embedding of the $j$-th candidate initially retrieved among the 100 documents. The reranked top-$k$ summaries are then incorporated into the prompts, enriching the contextual alignment of the generated responses. In detail, for each test individual, the retrieved and reranked summaries are combined into the final prompt, and the answer is subsequently generated using the function $f_{\text{gen}}$:
$P_{\text{test}}$ $=\left(D_{\text{test}},R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K},Q_{\text{test}}\right),$ (7)
$A_{\text{test}}$ $=f_{\text{gen}}\left(P_{\text{test}}\right).$
$P_{\text{test}}$ is the final prompt, $D_{\text{test}}$ is the demographic information of the test individual, $\{R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K}\}$ represents the top-$k$ reranked summaries, and $Q_{\text{test}}$ is the test question. $A_{\text{test}}$ denotes the generated answer for the test question, and $f_{\text{gen}}$ represents the generation function that leverages the structured input prompt.
This retrieval-based approach, followed by reranking, enhances reasoning by explicitly guiding the LLM to critically evaluate which retrieved values best align with the test individual’s demographic characteristics. The final prompts are then used to generate answers following the chain-of-thought prompting strategy, ensuring that the responses are contextually coherent and culturally aligned with the test individual’s profile.
5 Experiments
5.1 Setup
Models Used
We employ GPT-4o-mini Achiam et al. (2023), a smaller version of GPT-4o, for our generation tasks, with the temperature parameter set to $0.7$ for balancing coherence and creativity. For retrieval task, we used the E5 (base) model Wang et al. (2022) [^2] https://huggingface.co/intfloat/e5-base to generate embeddings and retrieve the top-100 most relevant summaries of values by calculating cosine similarity. Additionally, we utilized the GTE-multilingual-reranker-base model Zhang et al. (2024) [^3] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking. This reranking model refines the selection process and ensures that the top-k most relevant summaries are selected from the initial top-100 retrievals.
Baseline Methods and Implementation
Our baseline methods include: (1) Zero-shot Inference, (2) the role-assignment-only approach Tao et al. (2024), (3) a few-shot learning method Choenni and Shutova (2024), and (4) a hybrid method that combines both (1) and (2). Specifically, for the role-assignment baseline, we use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data. For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts. The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set.
Evaluation Method
We used accuracy as the primary evaluation metric, converting multi-choice responses into binary form for consistency and simplicity by dividing responses into two distinct groups based on contrasting answer patterns, which effectively captures agreement or disagreement and aligns with the structure of values-related questions.
5.2 Experimental Analysis
In our experiments, we compare our method ValuesRAG with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2. We found that the role-assignment method generally surpasses both zero-shot and few-shot approaches. By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance. In contrast, few-shot learning, which depends on a small set of examples, often faces challenges in generalizing to unseen scenarios, as different values dimensions are not strongly related in many cases. However, while role-assignment offers clearer cultural context, it may oversimplify complex personal traits by imposing rigid, stereotype-like constraints on agent behavior. The hybrid method seeks to address these limitations by combining role assignment for consistency with few-shot prompts to introduce variety, but still falls short in capturing the full complexity of cultural values.
In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific demographic and cultural data for each agent. Specifically, ValuesRAG (k=$3$) achieves the best results, with ValuesRAG (k=$5$) following closely behind. Although using $1$ or $10$ retrieved summaries also achieved better performance than other baselines, retrieving $3$ and $5$ summaries provides a more balanced trade-off between retrieval diversity and contextual relevance. This retrieval-augmented framework enables the model to dynamically incorporate richer, more nuanced information about each agent, avoiding the limitations of rigid demographic labels or small sample prompts. As a result, our model captures the interplay between individual beliefs, social contexts, and cultural norms more effectively. Evaluations across diverse test datasets demonstrate that ValuesRAG consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
5.3 Ablation Study: Impact of Values-Only Generation
To validate the robustness of ValuesRAG, we perform an ablation study using only value-augmented generation, thereby isolating the impact of demographic summaries on the model’s performance. Specifically, we used the WVS validation set—separated from the training data as outlined in Section 3—to evaluate the models. Table 3 presents a comparison of our method, using only summaries of values, against four baseline methods. The results demonstrate that ValuesRAG consistently outperforms the baselines across all datasets, achieving the highest accuracy even relying solely on summaries of values.
Methods Accuracy
Zero-Shot 0.6176
Role-Assignment 0.6747
Few-Shot Learning 0.6359
Role-Assignment + Few-Shot Learning 0.6670
Values Augmented Generation-Only (Ours) 0.6894
Table 3: Accuracy comparison between baseline methods and our proposed Values Augmented Generation method using the WVS validation set. Bold text indicates the best performance, underlined text indicates the second-best performance.
This result confirms the effectiveness and robustness of the values-augmented generation approach. ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses. Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined prompts or demographic anchors. The result showcases the framework’s scalability and adaptability, proving it can mitigate biases and generate culturally coherent outputs with minimal reliance on external context.
6 Conclusion
We present ValuesRAG, an innovative Retrieval-Augmented Generation framework that revolutionizes cultural values alignment through its comprehensive approach to values context analysis. While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles. Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines across diverse regional QA datasets. By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses. This structured design enables ValuesRAG to bridge the gap between generic LLM capabilities and the demands of culturally sensitive applications, providing a scalable and robust solution for real-world use cases.
7 Limitation
Although our baseline comparisons indicate that ValuesRAG generally delivers superior performance compared to alternative methods, it does not always guarantee an exact match to individual’s true values. Since we rely on the WVS dataset to summarize individual profiles, there can be mismatches when these summaries are applied to other test sets. In future work, we plan to explore more adaptive retrieval strategies that can better align with novel datasets, as well as investigate how integrating additional fine-tuning with retrieval-augmented generation may further refine each agent’s contextual accuracy.

### v2

@@
 Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively.
-To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with in-context learning to integrate cultural and demographic knowledge dynamically during text generation.
+To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with In-Context Learning (ICL) to integrate cultural and demographic knowledge dynamically during text generation.
 Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual.
-Subsequently, we curated several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries.
-ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided, highlighting ValuesRAG’s potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
+Subsequently, we curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries.
+ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided.
+Notably, ValuesRAG demonstrates an accuracy of 21% improvement over other baseline methods, highlighting its potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
 1 Introduction
@@
 While some countries have developed localized LLMs, such as China’s ERNIE Sun et al.
-(2021), South Korea’s HyperCLOVA Yoo et al.
-(2024), and the multilingual ChatGLM GLM et al.
+(2021), ChatGLM GLM et al.
+(2024), DeepSeek Liu et al.
+(2024a), and South Korea’s HyperCLOVA Yoo et al.
 (2024), these models also exhibit biases inherited from their respective training datasets.
@@
 (2024).
-Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and the proposed ValuesRAG framework.
-Recent studies have proposed several approaches, such as role-assignment approaches and few-shot learning techniques  Li et al.
-(2024a); AlKhamissi et al.
-(2024), to mitigate these cultural biases.
+Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and our proposed ValuesRAG framework.
+Recent studies have proposed several approaches, such as role-assignment approaches Tao et al.
+(2024) and few-shot learning techniques  Choenni and Shutova (2024), to mitigate these cultural biases.
 However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
-To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and in-context learning to dynamically incorporate cultural knowledge during text generation (see Figure 1).
+To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL) to dynamically incorporate cultural knowledge during text generation (see Figure 1).
 Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al.
@@
 After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected.
-Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies step-by-step reasoning grounded in the retrieved values to generate the final answer to the question.
+Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies reasoning grounded in the retrieved values to generate the final answer to the question.
 We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al.
@@
 To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs.
-Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines, also in the ablation study where only summaries of values were provided without any additional information.
-Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
+Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines.
+Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
+In addition, ablation studies on using only value-augmented generation and on varying the number of retrieved summaries confirm ValuesRAG’s robust performance under different configurations.
+ValuesRAG surpasses the baselines through purely values-based generation.
+Meanwhile, adjusting the number of retrieved documents shows that moderate retrievals can effectively balance diversity and relevance, also maintaining high accuracy across multiple benchmarks.
 These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications.
-Our study demonstrates ValuesRAG’s robust capabilities on a global scale, suggesting its applicability in aligning the values of diverse groups within a single country.
+Our study demonstrates ValuesRAG’s robust capabilities on a global scale, also suggesting its applicability in aligning the values of diverse groups within a single country.
 ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al.
-(2024b).
+(2024).
 This, in turn, facilitates the creation of fairer and more effective policies.
@@
 Unlike the more obvious safety concerns and social discrimination Liu et al.
-(2024) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data.
+(2024b) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data.
 Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al.
@@
 These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
-Set Dataset Abbreviation Region Year Na VQb
-Train Set World Values Survey WVS Global 2017–2022 97.2k 259
-Test Sets European Values Study EVS Europe 2017 59.4k 211
+Category Dataset Abbreviation Region Year N VQ
+Retrieval Corpus World Values Survey WVS Global 2017–2022 97.2k 259
+Test Datasets European Values Study EVS Europe 2017 59.4k 211
 The General Social Survey GSS North America 2021–2022 8.2k 44
@@
 Table 1: Overview of the datasets utilized in our study.
-The Train Set (WVS) includes global data collected between 2017 and 2022, providing the basis for generating cultural summaries of values and validation for our method.
-The Test Set consists of six region-specific surveys, each capturing socio-cultural information from distinct geographic areas and time frames.
-aN represents sample size in thousands (k).
-bVQ represents the number of values-related questions.
-cDQ represents the number of demographic features.
+The Retrieval Corpus (WVS) includes global data collected between 2017 and 2022, providing the basis for generating cultural summaries of values and validation for our method.
+The Test Datasets consist of six region-specific surveys, each capturing socio-cultural information from distinct geographic areas and time frames.
+N represents sample size in thousands (k).
+VQ represents the number of values-related questions.
 2.2 Mitigation of LLMs’ Cultural Bias
@@
 (2023); Ji et al.
-(2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al.
+(2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not that suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al.
 (2024), as there are hundreds of countries and cultures in the earth.
@@
 Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity.
-We, therefore, use these methods as baselines to benchmark our proposed approach.
+We therefore use these methods as baselines to benchmark our proposed approach.
 3 Datasets
-In this section, we first introduce the World Values Survey (WVS) as our training dataset (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies.
+In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies.
 Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
-3.1 Train Dataset
-WVS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al.
+3.1 Retrieval Corpus
+VS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al.
 (2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries.
-WVS is selected as our train dataset especially due to its numerous advantages:
+WVS is selected as our retrieval corpus especially due to its numerous advantages:
 1. Broad recognition and inclusiveness:
-WVS is widely recognized and used by governments, social scientists, and international organizations in values studies.
-It covers 120 countries, representing 94.5% of the global population, ensuring broad geographic and cultural representation.
+WVS is widely recognized and frequently used by governments, social scientists, and major international organizations in comparative values studies.
+It currently covers 120 countries, representing 94.5% of the global population, ensuring broad geographic and cultural representation.
 2. Expert-designed and accessible:
-The dataset is designed by leading domain experts to conduct comprehensive surveys of values, ensuring reliability and relevance.
+The dataset is meticulously designed by leading domain experts to conduct comprehensive surveys of values, ensuring reliability, rigor, and relevance.
 It is publicly accessible, enabling reproducibility and transparency in research.
@@
 WVS has well-organized and comprehensive demographic questions, making it effective for retrieval tasks. Its large sample size (97,221 respondents) is also suitable for RAG tasks.
-Since values evolve gradually over time, WVS is conducted in waves, with each wave occurring every five years. For our study, we utilized the most recent wave, spanning from 2017 to 2022.
+Since values evolve gradually over time, VS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
 The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions.
-The values questions span 13 topics, such as social trust, post-materialism, and political interest.
-We randomly selected 20% (52 questions) per topic for validation and used the remaining 80% (207 questions) for summary generation.
-The 31 demographic features, including country, sex, age, education, social class, and employment status, were used to generate demographic summaries for retrieval tasks.
+The value questions span 13 topics, such as social trust, post-materialism, and political interest.
+We randomly select 20% (52 questions) per topic for validation and use the remaining 80% (207 questions) for summary generation.
+The 31 demographic features, including country, sex, age, education, social class, and employment status, are used to generate demographic summaries for retrieval tasks.
+Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment.
+Consisting four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
 3.2 Test Datasets
-We selected six regional surveys to serve as test datasets based on the following criteria:
+We select six regional surveys [^2] Detailed dataset description are provided in Appendix A. to serve as test datasets based on the following criteria:
 1. Demographic and values coverage:
@@
 2. Temporal proximity:
-The datasets exhibit temporal proximity to WVS Wave 7 (2017–2022), allowing aligned comparisons and ensuring consistency across evaluations.
-The regions in our test datasets were meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring the data accurately reflects the majority of the global population.
-All of them are publicly accessible, and are statistically representative at national or regional levels, which guarantees their reliability and validity.
-To be specific, we selected European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region.
-For the United States, we selected the General Social Survey Davern et al.
+The datasets exhibit close temporal proximity to WVS Wave 7 (2017–2022), thereby allowing aligned comparisons and ensuring thorough consistency across diverse global evaluations.
+The regions in our test datasets are meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring that the data accurately reflects the majority of the global population.
+All of them are publicly accessible and are statistically representative at national or regional levels, which guarantees their reliability and validity.
+To be specific, we select the European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region.
+For the United States, we select the General Social Survey Davern et al.
 (2024), which is the most comprehensive social survey in the country.
 The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor.
-For India, where national survey data were largely inaccessible during our study, we used Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population.
-The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, was selected to represent Latin American countries, as it covers 32 countries across the region.
-Finally, Afrobarometer Afrobarometer (2023) was chosen as the representative dataset for Africa.
+For India, where national survey data were largely inaccessible during our study, we use Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population.
+The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, is selected to represent Latin American countries, as it covers 32 countries across the region.
+Finally, Afrobarometer Afrobarometer (2023) is chosen as the representative dataset for Africa.
 A detailed summary of these datasets is presented in Table 1.
-Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment, consisting of four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
 4 Methodology
@@
 ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3).
+[^3] We provide a case study of retrieve in Appendix C.
 An overview of the ValuesRAG framework is provided in Figure 2.
 4.1 Values and Demographic Summary Generation
-To systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages.
+systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages.
 First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1).
@@
 Finally, individual summaries are constructed by combining all topic summaries:
-$\small S_{i}=f_{\text{gen}}(T_{i}^{1},T_{i}^{2},\dots,T_{i}^{M}),$ (2)
+$S_{i}=f_{\text{gen}}(T_{i}^{1},T_{i}^{2},\dots,T_{i}^{M}),$ (2)
 with $M$ denoting the total number of topics.
@@
 4.2 Values Augmented Generation
-Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating the validation question-answer data.
+Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating on the validation question-answer data.
 For each validation question, we concatenate the corresponding individual’s values summary with the question itself, forming a context-rich input for the LLM:
@@
 To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval.
-The demographic data from both the train and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model.
+The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model.
 We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
-Methods EVS GSS CGSS ISD LAPOP Africa Avg.
-Baseline
-Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
-Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
-Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
-Role-Assignment + Few-Shot Learning 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
-Ours
-ValuesRAG($k=1$) 0.5960 0.7722* 0.5347* 0.6853* 0.4682 0.5904* 0.6078*
-ValuesRAG($k=3$) 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
-ValuesRAG($k=5$) 0.6051* 0.7706* 0.5301* 0.7016* 0.5061* 0.5905* 0.6173*
-ValuesRAG($k=10$) 0.6020* 0.7380 0.5317* 0.7014* 0.4686 0.5680* 0.6016*
-Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets, including the European Values Study (EVS), General Social Survey (GSS), Chinese General Social Survey (CGSS), Indian Survey Data (ISD), LAPOP, Afrobarometer, and the AmericasBarometer.
-$k$ indicates the number of summaries to be retrieved.
-Bold text indicates the best performance, underlined text indicates the second-best performance.
-* denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p$ <0.05) over all baseline model(s).
 $\small\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
-Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score.
+Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score.
 The top-100 embeddings with the highest similarity scores are initially selected as candidates.
@@
 $A_{\text{test}}$ $=f_{\text{gen}}\left(P_{\text{test}}\right).$
-$P_{\text{test}}$ is the final prompt, $D_{\text{test}}$ is the demographic information of the test individual, $\{R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K}\}$ represents the top-$k$ reranked summaries, and $Q_{\text{test}}$ is the test question.
-$A_{\text{test}}$ denotes the generated answer for the test question, and $f_{\text{gen}}$ represents the generation function that leverages the structured input prompt.
+Here, $P_{\text{test}}$ is the final prompt, $D_{\text{test}}$ is the demographic information of the test individual, $\{R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K}\}$ represents the top-$k$ reranked summaries, and $Q_{\text{test}}$ is the test question.
+$A_{\text{test}}$ denotes the generated answer for the test question, and $f_{\text{gen}}$ represents the generation function.
 This retrieval-based approach, followed by reranking, enhances reasoning by explicitly guiding the LLM to critically evaluate which retrieved values best align with the test individual’s demographic characteristics.
 The final prompts are then used to generate answers following the chain-of-thought prompting strategy, ensuring that the responses are contextually coherent and culturally aligned with the test individual’s profile.
+For the comprehensive implementation of the ValuesRAG, we provide the Algorithm 1, which systematically outlines the processes of values and demographic summary generation, values-augmented generation, and retrieval-based values alignment, as shown below:
+Algorithm 1 Values Generation and Retrieval Process
+1:Dataset $\mathcal{D}$ with topics and demographic QA pairs, Generative Model $f_{\text{gen}}$, Embedding Model $f_{\text{embed}}$, Reranking Model $f_{\text{rerank}}$, Retrieval Top-$K$
+2:// Values and Demographic Summary Generation
+3:for each individual $i\in\mathcal{D}$ do
+4:     Generate topic-based values summaries $T_{i}^{j}$ for each topic $j$
+5:     Generate demographic summaries $D_{i}$
+6:     Combine $T_{i}^{j}$ into comprehensive values summary $S_{i}$
+7:end for
+8:// Values Augmented Generation
+9:for each validation question $Q_{i}^{\text{val},k}$ of individual $i$ do
+10:     Construct context $C_{i}=\text{concat}(S_{i},Q_{i}^{\text{val},k})$
+11:     Augment context with demographic summary $D_{i}$
+12:     Generate answer $A_{i}=f_{\text{gen}}(C_{i},D_{i})$
+13:end for
+14:// Retrieval-based Values Alignment
+15:for each test individual $i$ do
+16:     Compute embeddings $E_{\text{test}}=f_{\text{embed}}(D_{\text{test}})$
+17:     Retrieve top-$100$ values summaries by similarity:
+18:      $\text{Sim}(E_{\text{test}},E_{\text{train}})$
+19:     Rerank top-$K$ summaries:
+20:      $R^{\prime}_{k}=f_{\text{rerank}}(E_{\text{test}},E_{C_{j}})$
+21:     Final prompt $P_{\text{test}}=(D_{\text{test}},R^{\prime}_{1},\dots,R^{\prime}_{K},Q_{\text{test}})$
+22:     Generate answer $A_{\text{test}}=f_{\text{gen}}(P_{\text{test}})$
+23:end for
+Model Methods EVS GSS CGSS ISD LAPOP Africa Avg.
+GPT-4o mini Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
+Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
+Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
+Hybrid Method 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
+ValuesRAG 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
+Gemini 1.5 Flash Zero-shot Inference 0.5419 0.6408 0.4502 0.6017 0.4149 0.4181 0.5113
+Role-Assignment 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
+Few-Shot Learning 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
+Hybrid Method 0.5845 0.7193 0.5026 0.6253 0.4448 0.5166 0.5655
+ValuesRAG 0.5869 0.7686* 0.5337* 0.6789* 0.4705 0.5473* 0.5977*
+Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets.
+$k$ indicates the number of summaries to be retrieved.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
+* denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p<0.05$) over all baseline model(s).
+denotes our proposed method.
+Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg.
+GPT-4o mini 1 0.5960 0.7722 0.5347 0.6853 0.4682 0.5905 0.6078
+3 0.6021 0.7781 0.5387 0.7001 0.4686 0.5953 0.6138
+5 0.6052 0.7706 0.5301 0.7016 0.5061 0.5905 0.6174
+10 0.6020 0.7380 0.5317 0.7014 0.5030 0.5680 0.6074
+Gemini 1.5 Flash 1 0.5753 0.7668 0.5272 0.6646 0.4548 0.5369 0.5876
+3 0.5869 0.7686 0.5337 0.6789 0.4705 0.5473 0.5977
+5 0.5868 0.7690 0.5303 0.6734 0.4661 0.5498 0.5959
+10 0.5852 0.7665 0.5279 0.6773 0.4509 0.5464 0.5924
+Table 3: Accuracy scores across six regional datasets for the ablation study (Section 5.3.2).
+Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
 5 Experiments
 5.1 Setup
-Models Used
-We employ GPT-4o-mini Achiam et al.
-(2023), a smaller version of GPT-4o, for our generation tasks, with the temperature parameter set to $0.7$ for balancing coherence and creativity.
-For retrieval task, we used the E5 (base) model Wang et al.
-(2022) [^2] https://huggingface.co/intfloat/e5-base to generate embeddings and retrieve the top-100 most relevant summaries of values by calculating cosine similarity.
-Additionally, we utilized the GTE-multilingual-reranker-base model Zhang et al.
+Models Used.
+We utilize GPT-4o-mini Achiam et al.
+(2023) and Gemini-1.5-Flash Team et al.
+(2024) for our generation tasks, which are accessed via APIs.
+We set the temperature parameter of these models to $0.7$ to achieve a balance between coherence and creativity.
+For the retrieval task, we employed the E5 (base) model Wang et al.
+(2022) [^4] https://huggingface.co/intfloat/e5-base, which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity.
+Additionally, we utilize the GTE-multilingual-reranker-base model Zhang et al.
 (2024) [^5] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking.
 This reranking model refines the selection process and ensures that the top-$k$ most relevant summaries are selected from the initial top-100 retrievals.
-Baseline Methods and Implementation
+Baseline Methods and Implementation.
 Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach Tao et al.
@@
 The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set.
-Evaluation Method
-We used accuracy as the primary evaluation metric, converting multi-choice responses into binary form for consistency and simplicity by dividing responses into two distinct groups based on contrasting answer patterns, which effectively captures agreement or disagreement and aligns with the structure of values-related questions.
+[^6] Detailed prompts are provided in Appendix B.
+Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance.
+[^7] Detailed analysis of varying $k$ and its implications on model performance is provided in Section 5.3.1.
+Evaluation.
+We utilize accuracy as the primary evaluation metric by converting multiple-choice responses into a binary format for consistency and simplicity.
+This is achieved by categorizing responses into two distinct groups based on contrasting answer patterns, effectively capturing agreement and disagreement.
+This approach aligns with the structure of values-related questions, such as those using a Likert Scale.
 5.2 Experimental Analysis
-In our experiments, we compare our method ValuesRAG with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2.
-We found that the role-assignment method generally surpasses both zero-shot and few-shot approaches.
+In our experiments, we compare our method, ValuesRAG, with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2.
+We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches.
 By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance.
-In contrast, few-shot learning, which depends on a small set of examples, often faces challenges in generalizing to unseen scenarios, as different values dimensions are not strongly related in many cases.
-However, while role-assignment offers clearer cultural context, it may oversimplify complex personal traits by imposing rigid, stereotype-like constraints on agent behavior.
-The hybrid method seeks to address these limitations by combining role assignment for consistency with few-shot prompts to introduce variety, but still falls short in capturing the full complexity of cultural values.
-In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific demographic and cultural data for each agent.
-Specifically, ValuesRAG (k=$3$) achieves the best results, with ValuesRAG (k=$5$) following closely behind.
-Although using $1$ or $10$ retrieved summaries also achieved better performance than other baselines, retrieving $3$ and $5$ summaries provides a more balanced trade-off between retrieval diversity and contextual relevance.
-This retrieval-augmented framework enables the model to dynamically incorporate richer, more nuanced information about each agent, avoiding the limitations of rigid demographic labels or small sample prompts.
-As a result, our model captures the interplay between individual beliefs, social contexts, and cultural norms more effectively.
-Evaluations across diverse test datasets demonstrate that ValuesRAG consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
-5.3 Ablation Study: Impact of Values-Only Generation
-To validate the robustness of ValuesRAG, we perform an ablation study using only value-augmented generation, thereby isolating the impact of demographic summaries on the model’s performance.
-Specifically, we used the WVS validation set—separated from the training data as outlined in Section 3—to evaluate the models.
+Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically.
+Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings.
+As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios.
+The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity.
+However, it remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
+In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific values-related cultural data for each agent.
+By focusing on values as the primary retrieval targets, this retrieval-augmented framework enables the model to include an expansive set of contextual clues, helping it reflect the depth and breadth of each individual’s background and values.
+Crucially, our ValuesRAG provides a more adaptive mechanism for representing the subtle interplay of personal beliefs and cultural norms via avoiding the limitations of rigid demographic labels or small-sample prompts.
+ValuesRAG more effectively captures the complex dynamics that can shape a respondent’s stance on different questions.
+Evaluations across diverse test datasets demonstrate that ValuesRAG with $k=3$ consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
+5.3 Ablation Study
+We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG., in Section 5.3.1, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance.
+Second, in Section 5.3.2, we isolate the effect of using only values-based generation.
+5.3.1 5.3.1 Varying the Number of Retrieved Summaries
+To further investigate the role of retrieval depth in ValuesRAG, we vary the number of retrieved summaries, $k\in\{1,3,5,10\}$.
+Table 3 presents the performance of ValuesRAG for three models across multiple datasets.
+We observe that while $k=1$ may not provide enough diversity, increasing $k$ beyond 3 can sometimes introduce less relevant information.
+Moreover, $k=3$ consistently achieves significantly better accuracy compared to higher values of $k$, highlighting its effectiveness and achieving the best balance between contextual relevance and model latency; these ensure both effectiveness and efficiency.
+This configuration not only enhances retrieval accuracy but also maintains low computational overhead, underscoring the strength of ValuesRAG in real-world applications where speed and performance are equally crucial.
+Hence, we choose $k=3$ as the default retrieval number for conducting our baseline experiments, ensuring optimal performance across various models and datasets.
+5.3.2 5.3.2 Impact of Values-Only Generation
+To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance.
+We use the WVS validation set—separated from the training data, which served as the retrieval corpus (as outlined in Section 3)—to evaluate the models.
 Table 4 presents a comparison of our method, using only summaries of values, against four baseline methods.
-The results demonstrate that ValuesRAG consistently outperforms the baselines across all datasets, achieving the highest accuracy even relying solely on summaries of values.
-Methods Accuracy
-Zero-Shot 0.6176
-Role-Assignment 0.6747
-Few-Shot Learning 0.6359
-Role-Assignment + Few-Shot Learning 0.6670
-Values Augmented Generation-Only (Ours) 0.6894
+Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
+Methods GPT-4o-mini Gemini-1.5-flash
+Zero-Shot 0.6176 0.6041
+Role-Assignment 0.6747 0.6505
+Few-Shot Learning 0.6359 0.6086
+Hybrid Method 0.6670 0.6354
+Values Augmented 0.6894 0.6583
 Table 4: Accuracy comparison between baseline methods and our proposed Values Augmented Generation method using the WVS validation set.
@@
 ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses.
-Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined prompts or demographic anchors.
-The result showcases the framework’s scalability and adaptability, proving it can mitigate biases and generate culturally coherent outputs with minimal reliance on external context.
+Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors.
+The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
 6 Conclusion
@@
 While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles.
-Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines across diverse regional QA datasets.
+Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines.
 By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses.
@@
 In future work, we plan to explore more adaptive retrieval strategies that can better align with novel datasets, as well as investigate how integrating additional fine-tuning with retrieval-augmented generation may further refine each agent’s contextual accuracy.
+Appendix A Appendix A. Dataset Description
+A.1 WVS
+As shown in Figure 3, the WVS dataset conducts surveys in over 120 countries worldwide, covering most regions and countries across the globe, making it suitable for constructing a documents corpus for Retrieval Augmented Generation (RAG).
+It is also representative in terms of demographic characteristics.
+We select the most recent WVS survey (Wave 7) to ensure the dataset is as close as possible to the contemporary world.
+Figure 3: Distribution of Country in WVS dataset. The WVS dataset is the most extensive values survey dataset in terms of coverage.
+According to the classification in the codebook released by WVS, the values-related questions it contains can be divided into 13 topics, as shown in Table 5, covering most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values.
+Topic Count
+Social Values, Norms, Stereotypes 45
+Happiness and Wellbeing 11
+Social Capital, Trust and Organizational Membership 47
+Economic Values 6
+Perceptions of Corruption 9
+Perceptions of Migration 10
+Perceptions of Security 21
+Index of Postmaterialism 6
+Perceptions about Science and Technology 6
+Religious Values 12
+Ethical Values 23
+Political Interest and Political Participation 35
+Political Culture and Political Regimes 25
+Table 5: Distribution of Values-related Questions in WVS. The questions were categorized into 13 topics with a total of 259 questions covering most of the dimensions of values
+A.2 Test Datasets
+To validate the robustness of our method, we selected datasets from six different regions as test sets.
+These test sets include both values-related questions and demographic characteristics.
+The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG.
+The values-related questions are utilized as test questions to calculate accuracy (ACC).
+EVS
+The first dataset comes from the European Values Study [^8] https://europeanvaluesstudy.eu, a large-scale, cross-national, and longitudinal survey research program designed to explore values, beliefs, and attitudes across Europe.
+This dataset includes a total of 211 values-related questions and captures 34 demographic characteristics of the respondents.
+We select EVS 2017, conducted in 2017, ensuring alignment with the World Values Survey (WVS) in terms of the timeframe.
+GSS
+We select GSS to represent the population of the United States.
+The General Social Survey [^9] https://gss.norc.org is a sociological survey that has been conducted since 1972 by the National Opinion Research Center (NORC) at the University of Chicago.
+Its primary purpose is to collect and analyze data on the opinions, behaviors, and demographic characteristics of adults in the United States, thereby monitoring societal change and the growing complexity of American society.
+Its questionnaire covers a comprehensive and wide range of topics, including many values-related questions.
+Specifically, within the GSS, we identify 44 questions as values-related and 33 questions as demographic characteristics.
+CGSS
+The Chinese General Social Survey [^10] http://cgss.ruc.edu.cn, initiated in 2003, is China’s earliest national, comprehensive, and continuous academic survey project.
+Conducted by the National Survey Research Center at Renmin University of China, the CGSS systematically collects data at multiple levels, including society, communities, families, and individuals.
+CGSS only provided the questionnaire and data in Chinese, which we have translated into English to ensure its usability.
+We ultimately compile a total of 58 values-related questions and 13 demographic characteristics.
+ISD
+To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set.
+However, we were unable to obtain data from several government surveys in India, so we used data published by the Pew Research Center instead.
+The Pew Research Center’s India Survey Dataset [^11] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance.
+Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others.
+This dataset covers 33 values-related questions and 23 demographic characteristics.
+LAPOP
+The Latin American Public Opinion Project (LAPOP) [^12] https://www.vanderbilt.edu/lapop is a research institute based at Vanderbilt University in Nashville, Tennessee.
+LAPOP’s most notable survey is the AmericasBarometer, the most extensive survey of democratic public opinion and behavior covering the Americas, including North, Central, South America, and the Caribbean.
+This survey measures democratic values and behaviors through voter surveys, providing valuable insights into public sentiments across the region.
+We select this dataset to represent the population of Latin America.
+There are 48 values-related questions and 12 demographic characteristics.
+Africa
+Afrobarometer [^13] https://www.afrobarometer.org is a pan-African, non-partisan research network established in 1999 that conducts public attitude surveys on democracy, governance, economic conditions, and related issues across Africa.
+We selected data from the 8th round of Afrobarometer (collected in 2022), covering 34 African countries.
+After screening, a total of 144 values-related questions and 14 demographic characteristics are obtained.
+Appendix B Appendix B. Prompts Used
+This section provides the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries.
+Prompt for Question Answering Task: Respond to the question as the target individual, selecting the answer that aligns with their values and demographic context.
+Rules: Step-by-step analysis using retrieved demographics and values data.
+Maintain the target individual’s perspective throughout the analysis.
+Provide the response in JSON format, without additional explanation.
+Steps for Inferring: 1.
+Analyze the demographics (age, gender, cultural background, social class, religion, and economic class) of retrieved individuals.
+Compare them with the target individual.
+2.
+Identify individuals whose demographics most closely match the target individual.
+Note their IDs.
+3.
+Based on the matched individual’s values, infer how the target would respond.
+4.
+Select the response that best aligns with the inferred values, and return only the integer representing the selected option.
+Prompt for Values Summary Generation You are a summarization expert with expertise in extracting key insights from complex data.
+Based on the provided context, summarize this person’s values in one paragraph.
+Prompt for Demographic Summary Generation You are a summarization expert with expertise in extracting key insights from complex data.
+Based on the provided context, summarize this person’s demographics in one paragraph.
+Appendix C Appendix C. Case Study
+Example 1 Question Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all?
+Original Demographics Summary The respondent is a 63-year-old male born in the United States, of White descent, who speaks English at home.
+He completed 12th grade and holds a high school diploma, while his spouse has the same educational background.
+They are currently married and live in a household of two, with no children.
+The respondent identifies as middle class and reported a family income of \$90,000 to \$109,999 last year, with his own earnings falling within the same range.
+He practices Protestantism and is affiliated with the United Presbyterian Church in the U.S.A.
+His mother has completed two years of college, while his father finished high school.
+The respondent’s demographic background reflects a stable family life with no financial challenges reported.
+Retrieved Summaries: Retrieved Text 1
+Demographics
+The respondent is a 60-year-old White, non-Hispanic male born in the United States, where he speaks English at home.
+He was raised in a family where both parents were also born in the U.S., and he holds a bachelor’s degree, signaling a level of educational attainment typical of the upper middle class.
+He is currently employed full-time in a professional and technical role within the private business sector and is the chief wage earner in his household, which consists of one person.
+The respondent is divorced and has two children.
+He identifies as a Protestant and has reported saving money over the past year, rating his household income as relatively high on a scale of one to ten.
+Values
+This person’s values center on family, personal empowerment, and social equity, emphasizing a balanced life encompassing work, political engagement, and community well-being.
+They prioritize financial responsibility, health, and individual liberties while advocating equality and rejecting traditional gender roles.
+They approach trust cautiously, emphasize accountability in governance, and hold nuanced views on immigration and moral integrity.
+They believe in science and technology’s benefits and maintain personal faith amid diverse beliefs.
+Politically active, they support democratic principles and civil rights while voicing dissatisfaction with current institutions.
+Overall, they blend progressive values focused on social justice, environmental sustainability, and the power of citizen engagement in politics and public life.
+Retrieved Text 2
+Demographics
+The respondent is a 63-year-old White, non-Hispanic male born in the United States, where he speaks English at home.
+He is a citizen of the U.S., with both parents also being born in the country.
+He is married and lives in a household of two people, including himself and his spouse, who is retired.
+He has two children and both he and his spouse completed 12th grade, earning a high school diploma.
+He works full-time in a semi-skilled occupation within the private sector and identifies as belonging to the middle class, reporting that his household income places them in the third income tier.
+The respondent aligns with the Protestant religion and describes their financial situation as stable.
+Values
+This individual values family and work as life’s center, while moderately engaging in politics.
+They emphasize independence, hard work, tolerance, and determination in children, embrace progressive stances on gender and social issues, and place little importance on religious conformity.
+Their trust is selective, relying heavily on family, showing moderate faith in others, and opposing corruption, while valuing economic equality and environmental concerns.
+They uphold science and technology, favor personal freedom over security, and maintain moral integrity and accountability.
+Strongly engaged in civic matters, they are committed to democracy, civil rights, and active political participation, yet remain cautious about government roles.
+Overall, their values balance familial dedication, personal agency, ethical living, and a progressive yet pragmatic outlook on social and political issues.
+Retrieved Text 3
+Demographics
+The respondent is a 65-year-old white male, born in the United States, …
+Example 2 Question How many children to have is a personal matter, and the government should not interfere.
+Do you agree?
+Original Demographics Summary The respondent is a 27-year-old male from China of Han ethnicity, who speaks Chinese at home.
+He holds a postgraduate degree and has never engaged in any work for economic income.
+Currently, he lives alone without any additional household members and identifies as belonging to the middle class, describing his household’s economic status as average in his local area.
+He does not practice any religion and reports no personal income in the past year.
+Retrieved Summaries: Retrieved Text 1
+Demographics
+The respondent is a 28-year-old male of Chinese ethnicity, born in the country where he currently resides.
+He speaks Standard Chinese (Mandarin/Putonghua/Guoyu) at home and identifies as part of the working class.
+He has attained a Bachelor’s degree and is employed full-time in a clerical role within the private sector.
+He lives in a household of four, including his own parents, and is currently married but has no children.
+Both he and his spouse have a similar educational background and are employed full-time.
+His family’s financial situation indicates that they have spent some savings and borrowed money over the past year.
+He does not belong to any religious denomination and has reported his household income as being in the third income level out of ten.
+Values
+This individual holds a complex set of values that place a significant emphasis on family, work, and personal responsibility, while also promoting income equality and valuing hard work as a means of success.
+They have a moderate stance on political engagement, expressing skepticism toward certain institutions while maintaining high trust in organizations like the police and government.
+Their views reflect a critical perspective on issues such as immigration and gender roles, where they value cultural diversity but prioritize native citizens for job opportunities.
+They exhibit a commitment to environmental protection, scientific advancements, and ethical living, while acknowledging their non-religious identity.
+Overall, their values convey a blend of personal autonomy, accountability, and a pragmatic view on governance, demonstrating a preference for democratic principles intertwined with the belief in the importance of social order and national security.
+Retrieved Text 2
+Demographics
+The respondent is a 29-year-old male of Chinese descent, and speaks other Chinese dialects at home.
+He has attained a bachelor’s degree and works full-time in the sales sector within a private business or industry.
+Living alone, he is single and has no children.
+His family background includes parents with upper secondary education.
+Socially, he identifies with the working class and describes his household’s financial status as just getting by, placing his household income in the fourth income group on a scale of ten.
+He does not identify with any religious denomination.
+Values
+This person exhibits a set of values that emphasize the importance of family, hard work, and personal responsibility, reflecting a traditional view of societal roles and a commitment to instilling these qualities in children.
+They hold a pragmatic and somewhat secular worldview, expressing low interest in politics and religion while maintaining a positive attitude towards health, personal agency, and economic stability.
+They advocate for income equality and environmental sustainability, demonstrating concern for social equity and opposing corruption, while also valuing trust primarily within familial ties.
+Additionally, their perspective on immigrants is cautiously appreciative, and they prioritize freedom and economic progress, alongside a recognition of the complexities of modern ethical and moral issues.
+Overall, this individual’s values highlight a balance between familial loyalty, economic and environmental responsibilities, personal freedom, and a discerning approach to broader societal structures.
+Retrieved Text 3
+Demographics
+The respondent is a 25-year-old male of Chinese ethnicity, who speaks Standard Chinese languages, …

### v3

@@
 Abstract
-Cultural values alignment in Large Language Models (LLMs) is a critical challenge due to their tendency to embed Western-centric biases from training data, leading to misrepresentations and fairness issues in cross-cultural contexts.
-Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively.
+Ensuring cultural values alignment in Large Language Models (LLMs) remains a critical challenge, as these models often embed Western-centric biases from their training data, leading to misrepresentations and fairness concerns in cross-cultural applications.
+Existing approaches—such as role assignment and few-shot learning—struggle to address these limitations effectively due to their reliance on pre-trained knowledge, limited scalability, and inability to capture nuanced cultural values.
 To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with In-Context Learning (ICL) to integrate cultural and demographic knowledge dynamically during text generation.
 Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual.
-Subsequently, we curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries.
-ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided.
-Notably, ValuesRAG demonstrates an accuracy of 21% improvement over other baseline methods, highlighting its potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
-1 Introduction
-The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  Singh et al.
-(2024); Kharchenko et al.
-(2024); Hu et al.
-(2024).
-Predominantly trained on Western data sources Achiam et al.
-(2023); Touvron et al.
-(2023); Jiang et al.
-(2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts.
-These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs Gallegos et al.
-(2024); Xie et al.
-(2024); Potter et al.
-(2024); Huang et al.
-(2024).
+We subsequently curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries.
+We evaluate ValuesRAG using $6$ diverse regional datasets and show that it consistently outperforms baselines: including zero-shot, role-assignment, few-shot, and hybrid methods, both in main experiments and ablation settings.
+Notably, ValuesRAG achieves the best overall performance over prior methods, demonstrating its effectiveness in fostering culturally aligned and inclusive AI systems.
+Our findings underscore the potential of dynamic retrieval-based methods to bridge the gap between global LLM capabilities and localized cultural values.
+Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and our proposed ValuesRAG framework.
+Introduction
+The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  (Singh et al.
+2024; Kharchenko et al.
+2024; Hu et al.
+2024).
+Predominantly trained on Western data sources (Achiam et al.
+2023; Touvron et al.
+2023; Jiang et al.
+2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts.
+These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs (Gallegos et al.
+2024; Xie et al.
+2024; Potter et al.
+2024; Huang, Papyshev, and Wong 2024).
 Despite ongoing efforts to address these issues, existing strategies often fall short.
-While some countries have developed localized LLMs, such as China’s ERNIE Sun et al.
-(2021), ChatGLM GLM et al.
-(2024), DeepSeek Liu et al.
-(2024a), and South Korea’s HyperCLOVA Yoo et al.
-(2024), these models also exhibit biases inherited from their respective training datasets.
-As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment Gallegos et al.
-(2024); Xie et al.
-(2024); Potter et al.
-(2024).
-Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and our proposed ValuesRAG framework.
-Recent studies have proposed several approaches, such as role-assignment approaches Tao et al.
-(2024) and few-shot learning techniques  Choenni and Shutova (2024), to mitigate these cultural biases.
+While some countries have developed localized LLMs, such as China’s ERNIE (Sun et al.
+2021), ChatGLM (GLM et al.
+2024), DeepSeek (Liu et al.
+2024a), and South Korea’s HyperCLOVA (Yoo et al.
+2024), these models also exhibit biases inherited from their respective training datasets.
+As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment (Gallegos et al.
+2024; Xie et al.
+2024; Potter et al.
+2024).
+Recent studies have proposed several approaches, such as role-assignment approaches (Tao et al.
+2024) and few-shot learning techniques  (Choenni and Shutova 2024), to mitigate these cultural biases.
 However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
 To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL) to dynamically incorporate cultural knowledge during text generation (see Figure 1).
-Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al.
-(2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies.
-Specifically, we first generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel.
-After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected.
+Our framework leverages the World Values Survey (WVS) dataset (Haerpfer et al.
+2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies.
+First, we specifically generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel.
+After constructing the knowledge base, we retrieve the top-$100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected.
 Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies reasoning grounded in the retrieved values to generate the final answer to the question.
-We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al.
-(2024), (3) few-shot learning Choenni and Shutova (2024), and (4) a hybrid method combining (1) and (2).
+We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method (Tao et al.
+2024), (3) few-shot learning (Choenni and Shutova 2024), and (4) a hybrid method combining both (1) and (2).
 To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs.
 Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines.
-Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
-In addition, ablation studies on using only value-augmented generation and on varying the number of retrieved summaries confirm ValuesRAG’s robust performance under different configurations.
-ValuesRAG surpasses the baselines through purely values-based generation.
-Meanwhile, adjusting the number of retrieved documents shows that moderate retrievals can effectively balance diversity and relevance, also maintaining high accuracy across multiple benchmarks.
+Compared to previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
+In addition, ablation studies on varying the number of retrieved summaries and on using only value-augmented generation confirm ValuesRAG’s robust performance under different configurations.
+Adjusting the number of retrieved documents shows that moderate retrievals can balance diversity and relevance, also maintaining high accuracy across multiple benchmarks.
+Meanwhile, ValuesRAG surpasses the baselines through purely values-based generation.
 These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications.
 Our study demonstrates ValuesRAG’s robust capabilities on a global scale, also suggesting its applicability in aligning the values of diverse groups within a single country.
-ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al.
-(2024).
+ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes (Li, Gong, and Jiang 2024).
 This, in turn, facilitates the creation of fairer and more effective policies.
@@
 This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
-2 Related Work
-2.1 Evaluation of LLMs’ Cultural Bias
-Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue Tao et al.
-(2024).
-Unlike the more obvious safety concerns and social discrimination Liu et al.
-(2024b) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data.
-Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al.
-(2024); Durmus et al.
-(2024).
-This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  Manvi et al.
-(2024).
-Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models Gallegos et al.
-(2024).
-For example, Webster et al.
-Webster et al.
-(2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models, while Caliskan et al.
-Caliskan et al.
-(2017) pioneered the use of word embeddings as quantitative measures of bias.
+Related Work
+Evaluation of LLMs’ Cultural Bias
+Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue (Tao et al.
+2024).
+While safety concerns and social discrimination in language models are typically explicit and well-recognized (Liu et al.
+2024b), cultural biases often manifest subtly, reflecting dominant cultural perspectives embedded within training data.
+Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  (Manvi et al.
+2024; Durmus et al.
+2024).
+This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  (Manvi et al.
+2024).
+Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models (Gallegos et al.
+2024).
+For instance, Caliskan et al.
+(2017) pioneered the use of word embeddings as quantitative measures of bias, while Webster et al.
+(2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models.
 More recently, Karinshak et al.
-Karinshak et al.
 (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol.
@@
 VQ represents the number of values-related questions.
-2.2 Mitigation of LLMs’ Cultural Bias
-Methods like RLHF Shen et al.
-(2023); Ji et al.
-(2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not that suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al.
-(2024), as there are hundreds of countries and cultures in the earth.
+Mitigation of LLMs’ Cultural Bias
+Techniques such as Reinforcement Learning from Human Feedback (RLHF) (Shen et al.
+2023; Ji et al.
+2024) are commonly employed for aligning LLMs with human values.
+However, these single-dimensional alignment methods are insufficient for mitigating cultural bias, as cultural values are inherently diverse, dynamic, and context-dependent, varying significantly across different regions and societies (Huang, Papyshev, and Wong 2024).
 Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs.
 For instance, Tao et al.
-Tao et al.
 (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al.
-Masoud et al.
 (2024) developed a soft prompt tuning approach to mitigate bias.
-Also, Choenni and Shutova Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts.
+Moreover, Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts.
 However, these approaches face significant limitations in fully capturing the complexity of cultural alignment.
-Tao et al.’s technique mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations.
-Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity.
+Tao et al.’s approach (Tao et al.
+2024) mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations.
+Meanwhile, Choenni and Shutova’s few-shot learning approach (Choenni and Shutova 2024) similarly falls short of modeling cultural alignment in all its complexity.
 We therefore use these methods as baselines to benchmark our proposed approach.
-3 Datasets
-In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies.
-Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
-3.1 Retrieval Corpus
-VS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al.
-(2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries.
+Datasets
+In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus, highlighting its extensive coverage, global representativeness, and relevance for values-related studies.
+Subsequently, we describe six regional test datasets, which are carefully selected to ensure geographic, cultural, and demographic diversity.
+Retrieval Corpus
+WVS [^1] https://www.worldvaluessurvey.org/wvs.jsp (Haerpfer et al.
+2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries.
 WVS is selected as our retrieval corpus especially due to its numerous advantages:
@@
 WVS has well-organized and comprehensive demographic questions, making it effective for retrieval tasks. Its large sample size (97,221 respondents) is also suitable for RAG tasks.
-Since values evolve gradually over time, VS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
+Since values evolve gradually over time, WVS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
 The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions.
-The value questions span 13 topics, such as social trust, post-materialism, and political interest.
+The value questions span 13 topics, such as social trust, post-materialism, and political interest, as shown in Table 2.
+It covers most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values.
 We randomly select 20% (52 questions) per topic for validation and use the remaining 80% (207 questions) for summary generation.
 The 31 demographic features, including country, sex, age, education, social class, and employment status, are used to generate demographic summaries for retrieval tasks.
-Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment.
-Consisting four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
-3.2 Test Datasets
-We select six regional surveys [^2] Detailed dataset description are provided in Appendix A. to serve as test datasets based on the following criteria:
+Figure 2: Distribution of countries in the WVS dataset.
+The WVS dataset covers surveys conducted in over $120$ countries across all major regions, providing broad geographic and demographic representation for building a reliable retrieval corpus in RAG-based frameworks.
+Topic Count
+Social Values, Norms, Stereotypes 45
+Happiness and Wellbeing 11
+Social Capital, Trust and Organizational Membership 47
+Economic Values 6
+Perceptions of Corruption 9
+Perceptions of Migration 10
+Perceptions of Security 21
+Index of Postmaterialism 6
+Perceptions about Science and Technology 6
+Religious Values 12
+Ethical Values 23
+Political Interest and Political Participation 35
+Political Culture and Political Regimes 25
+Table 2: Distribution of Values-related Questions in WVS. The questions were categorized into 13 topics with a total of 259 questions covering most of the dimensions of values
+Test Datasets
+We select six regional surveys to serve as test datasets based on the following criteria:
 1. Demographic and values coverage:
@@
 All of them are publicly accessible and are statistically representative at national or regional levels, which guarantees their reliability and validity.
-To be specific, we select the European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region.
-For the United States, we select the General Social Survey Davern et al.
-(2024), which is the most comprehensive social survey in the country.
-The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor.
-For India, where national survey data were largely inaccessible during our study, we use Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population.
-The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, is selected to represent Latin American countries, as it covers 32 countries across the region.
-Finally, Afrobarometer Afrobarometer (2023) is chosen as the representative dataset for Africa.
-A detailed summary of these datasets is presented in Table 1.
-4 Methodology
+Also, these test sets include both values-related questions and demographic characteristics.
+The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG.
+The values-related questions are utilized as test questions to calculate accuracy (ACC).
+[^2] A detailed description of the method for computing accuracy (ACC) is provided in Section Experiments - Evaluation Metrics.
+The specific datasets used in our evaluation, along with their characteristics, are described below:
+EVS
+The first dataset comes from the European Values Study [^3] https://europeanvaluesstudy.eu, a large-scale, cross-national, and longitudinal survey research program designed to explore values, beliefs, and attitudes across Europe.
+This dataset includes a total of 211 values-related questions and captures 34 demographic characteristics of the respondents.
+We select EVS 2017, conducted in 2017, ensuring alignment with the World Values Survey (WVS) in terms of the timeframe.
+GSS
+We select GSS to represent the population of the United States.
+The General Social Survey [^4] https://gss.norc.org is a sociological survey that has been conducted since 1972 by the National Opinion Research Center (NORC) at the University of Chicago.
+Its primary purpose is to collect and analyze data on the opinions, behaviors, and demographic characteristics of adults in the United States, thereby monitoring societal change and the growing complexity of American society.
+Its questionnaire covers a comprehensive and wide range of topics, including many values-related questions.
+Specifically, within the GSS, we identify 44 questions as values-related and 33 questions as demographic characteristics.
+CGSS
+The Chinese General Social Survey [^5] http://cgss.ruc.edu.cn, initiated in 2003, is China’s earliest national, comprehensive, and continuous academic survey project.
+Conducted by the National Survey Research Center at Renmin University of China, the CGSS systematically collects data at multiple levels, including society, communities, families, and individuals.
+CGSS only provided the questionnaire and data in Chinese, which we have translated into English to ensure its usability.
+We ultimately compile a total of 58 values-related questions and 13 demographic characteristics.
+ISD
+To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set.
+However, we were unable to obtain data from several government surveys in India, thus we used data published by the Pew Research Center instead.
+The Pew Research Center’s India Survey Dataset [^6] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance.
+Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others.
+This dataset covers 33 values-related questions and 23 demographic characteristics.
+LAPOP
+The Latin American Public Opinion Project (LAPOP) [^7] https://www.vanderbilt.edu/lapop is a research institute based at Vanderbilt University in Nashville, Tennessee.
+LAPOP’s most notable survey is the AmericasBarometer, the most extensive survey of democratic public opinion and behavior covering the Americas, including North, Central, South America, and the Caribbean.
+This survey measures democratic values and behaviors through voter surveys, providing valuable insights into public sentiments across the region.
+We select this dataset to represent the population of Latin America.
+There are 48 values-related questions and 12 demographic characteristics.
+Africa
+Afrobarometer [^8] https://www.afrobarometer.org is a pan-African, non-partisan research network established in 1999 that conducts public attitude surveys on democracy, governance, economic conditions, and related issues across Africa.
+We selected data from the 8th round of Afrobarometer (collected in 2022), covering 34 African countries.
+After screening, a total of 144 values-related questions and 14 demographic characteristics are obtained.
+Figure 3: Overview of the proposed ValuesRAG framework for cultural alignment.
+The framework comprises four key stages: (1) Dataset preprocessing to separate training and test QA pairs from WVS and regional datasets, (2) Topic-wise summary generation using LLMs for each individual, (3) Aggregation of topic summaries into comprehensive individual values profiles, and (4) Retrieval-Augmented Generation that retrieves and reranks relevant value summaries based on demographic similarity to guide final response generation.
+Methodology
 In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach.
-ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3).
-[^3] We provide a case study of retrieve in Appendix C.
+ValuesRAG consists of three key components: (1) Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets; (2) Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context; and (3) Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles.
 An overview of the ValuesRAG framework is provided in Figure 3.
-4.1 Values and Demographic Summary Generation
-systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages.
-First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1).
+Values and Demographic Summary Generation
+To systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages.
+First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved, as described in previous section.
 In parallel, topic-based summaries and demographic summaries are generated separately.
@@
 The result, denoted as $S_{i}$, forms a comprehensive values summary for individual $i$.
-These generate summaries serve as structured references for retrieval in later stages and are also used to augment the validation set for evaluation.
-4.2 Values Augmented Generation
+These generate summaries serve as structured references for retrieval in later stages and are used to augment the validation set for evaluation as well.
+Values Augmented Generation
 Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating on the validation question-answer data.
 For each validation question, we concatenate the corresponding individual’s values summary with the question itself, forming a context-rich input for the LLM:
-$C_{i}=\text{concat}\left(S_{i},\text{Q}_{i}^{\text{val},k}\right)$ (3)
+$C_{i}=\text{concatenate}\left(S_{i},\text{Q}_{i}^{\text{val},k}\right)$ (3)
 where $C_{i}$ represents the combined context, $S_{i}$ is the values summary for individual $i$, and $\text{Q}_{i}^{\text{val},k}$ is the $k$-th validation question.
-Next, we concatenate $C_{i}$ with the demographic summary $D_{i}$ to further enhance the context, enabling the generation of responses based on both values and demographic information:
+We subsequently concatenate $C_{i}$ with the demographic summary $D_{i}$ to further enhance the context, enabling the generation of responses based on both values and demographic information:
 $A_{i}=f_{\text{gen}}\left(C_{i},D_{i}\right)$ (4)
 Here, $A_{i}$ represents the answer generated by the function $f$, and $(C_{i},D_{i})$ embeds the augmented context $C_{i}$ and demographic information $D_{i}$ into a structured input format.
-Additionally, we utilize chain-of-thought prompting to enhance reasoning and emulate the behavior of the corresponding individual, ensuring responses that are contextually aligned with the values captured in the summaries and demographic characteristics.
-4.3 Retrieval-based Values Alignment
+Additionally, we utilize Chain-of-Thought (CoT) prompting (Wei et al.
+2022) to enhance reasoning and emulate the behavior of the corresponding individual, ensuring responses that are contextually aligned with the values captured in the summaries and demographic characteristics.
+Retrieval-based Values Alignment
 To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval.
-The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model.
+The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a sentence-transformer-based model (Reimers and Gurevych 2019).
+[^9] A detailed description of the model used is provided in  Section Experiments - Models Used.
 We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
-$\small\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
-Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score.
+$\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
+$E_{\text{test}}$ and $E_{\text{train}}$ specifically represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score.
 The top-100 embeddings with the highest similarity scores are initially selected as candidates.
-Subsequently, we apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates.
-The reranking process evaluates the semantic relevance of the initial candidates based on a scoring function:
-$\small R^{\prime}_{k}=f_{\text{rerank}}\left(E_{\text{test}},E_{C_{j}}\right),\;C_{j}\in\{C_{1},C_{2},\dots,C_{100}\},\;k\leq 100$ (6)
-where $R^{\prime}_{k}$ represents the reranked summary for the $k$-th candidate selected from the final top-$k$ results, $f_{\text{rerank}}$ is the reranking function, and $E_{C_{j}}$ denotes the embedding of the $j$-th candidate initially retrieved among the 100 documents.
-The reranked top-$k$ summaries are then incorporated into the prompts, enriching the contextual alignment of the generated responses.
+We subsequently apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates.
+The reranking process evaluates the semantic relevance by passing each candidate summary embedding $E_{C_{j}}$ along with the test individual’s embedding $E_{\text{test}}$ through a neural reranker $f_{\text{rerank}}$, which outputs a relevance score $s_{j}$:
+$s_{j}=f_{\text{rerank}}\left(E_{\text{test}},E_{C_{j}}\right),\quad j\in\{1,2,\dots,100\}$ (6)
+We then sort the candidate summaries based on these scores in descending order and select the top-$k$ summaries with the highest scores as the final reranked set $R^{\prime}_{k}$:
+$R^{\prime}_{k}=\text{Top-}k\left(\{C_{j}\}_{j=1}^{100},\{s_{j}\}_{j=1}^{100}\right)$ (7)
+The reranked top-$k$ summaries $R^{\prime}_{k}$ are incorporated into the prompts, enriching the contextual alignment of the generated responses.
 In detail, for each test individual, the retrieved and reranked summaries are combined into the final prompt, and the answer is subsequently generated using the function $f_{\text{gen}}$:
@@
 2:// Values and Demographic Summary Generation
-3:for each individual $i\in\mathcal{D}$ do
+3:for each individual $i$ in $\mathcal{D}$ do
 4:     Generate topic-based values summaries $T_{i}^{j}$ for each topic $j$
@@
 23:end for
-Model Methods EVS GSS CGSS ISD LAPOP Africa Avg.
-GPT-4o mini Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
-Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
-Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
-Hybrid Method 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
-ValuesRAG 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
-Gemini 1.5 Flash Zero-shot Inference 0.5419 0.6408 0.4502 0.6017 0.4149 0.4181 0.5113
-Role-Assignment 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
-Few-Shot Learning 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
-Hybrid Method 0.5845 0.7193 0.5026 0.6253 0.4448 0.5166 0.5655
-ValuesRAG 0.5869 0.7686* 0.5337* 0.6789* 0.4705 0.5473* 0.5977*
-Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets.
-$k$ indicates the number of summaries to be retrieved.
-Bold text indicates the best performance, underlined text indicates the second-best performance.
-* denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p<0.05$) over all baseline model(s).
-denotes our proposed method.
-Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg.
-GPT-4o mini 1 0.5960 0.7722 0.5347 0.6853 0.4682 0.5905 0.6078
-3 0.6021 0.7781 0.5387 0.7001 0.4686 0.5953 0.6138
-5 0.6052 0.7706 0.5301 0.7016 0.5061 0.5905 0.6174
-10 0.6020 0.7380 0.5317 0.7014 0.5030 0.5680 0.6074
-Gemini 1.5 Flash 1 0.5753 0.7668 0.5272 0.6646 0.4548 0.5369 0.5876
-3 0.5869 0.7686 0.5337 0.6789 0.4705 0.5473 0.5977
-5 0.5868 0.7690 0.5303 0.6734 0.4661 0.5498 0.5959
-10 0.5852 0.7665 0.5279 0.6773 0.4509 0.5464 0.5924
-Table 3: Accuracy scores across six regional datasets for the ablation study (Section 5.3.2).
-Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved.
-Bold text indicates the best performance, underlined text indicates the second-best performance.
-5 Experiments
-5.1 Setup
+Experiments
+Setup
 Models Used.
-We utilize GPT-4o-mini Achiam et al.
-(2023) and Gemini-1.5-Flash Team et al.
-(2024) for our generation tasks, which are accessed via APIs.
+We utilize GPT-4o-mini (Achiam et al.
+2023) and Gemini-1.5-Flash (Team et al.
+2024) for our generation tasks, which are accessed via APIs.
 We set the temperature parameter of these models to $0.7$ to achieve a balance between coherence and creativity.
-For the retrieval task, we employed the E5 (base) model Wang et al.
-(2022) [^4] https://huggingface.co/intfloat/e5-base, which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity.
-Additionally, we utilize the GTE-multilingual-reranker-base model Zhang et al.
-(2024) [^5] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking.
-This reranking model refines the selection process and ensures that the top-$k$ most relevant summaries are selected from the initial top-100 retrievals.
-Baseline Methods and Implementation.
-Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach Tao et al.
-(2024), (3) a few-shot learning method Choenni and Shutova (2024), and (4) a hybrid method that combines both (1) and (2).
-Specifically, for the role-assignment baseline, we use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data.
-For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts.
-The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set.
-[^6] Detailed prompts are provided in Appendix B.
-Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance.
-[^7] Detailed analysis of varying $k$ and its implications on model performance is provided in Section 5.3.1.
-Evaluation.
-We utilize accuracy as the primary evaluation metric by converting multiple-choice responses into a binary format for consistency and simplicity.
-This is achieved by categorizing responses into two distinct groups based on contrasting answer patterns, effectively capturing agreement and disagreement.
-This approach aligns with the structure of values-related questions, such as those using a Likert Scale.
-5.2 Experimental Analysis
-In our experiments, we compare our method, ValuesRAG, with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2.
-We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches.
-By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance.
-Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically.
-Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings.
-As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios.
-The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity.
-However, it remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
-In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific values-related cultural data for each agent.
-By focusing on values as the primary retrieval targets, this retrieval-augmented framework enables the model to include an expansive set of contextual clues, helping it reflect the depth and breadth of each individual’s background and values.
-Crucially, our ValuesRAG provides a more adaptive mechanism for representing the subtle interplay of personal beliefs and cultural norms via avoiding the limitations of rigid demographic labels or small-sample prompts.
-ValuesRAG more effectively captures the complex dynamics that can shape a respondent’s stance on different questions.
-Evaluations across diverse test datasets demonstrate that ValuesRAG with $k=3$ consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
-5.3 Ablation Study
-We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG., in Section 5.3.1, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance.
-Second, in Section 5.3.2, we isolate the effect of using only values-based generation.
-5.3.1 5.3.1 Varying the Number of Retrieved Summaries
-To further investigate the role of retrieval depth in ValuesRAG, we vary the number of retrieved summaries, $k\in\{1,3,5,10\}$.
-Table 3 presents the performance of ValuesRAG for three models across multiple datasets.
-We observe that while $k=1$ may not provide enough diversity, increasing $k$ beyond 3 can sometimes introduce less relevant information.
-Moreover, $k=3$ consistently achieves significantly better accuracy compared to higher values of $k$, highlighting its effectiveness and achieving the best balance between contextual relevance and model latency; these ensure both effectiveness and efficiency.
-This configuration not only enhances retrieval accuracy but also maintains low computational overhead, underscoring the strength of ValuesRAG in real-world applications where speed and performance are equally crucial.
-Hence, we choose $k=3$ as the default retrieval number for conducting our baseline experiments, ensuring optimal performance across various models and datasets.
-5.3.2 5.3.2 Impact of Values-Only Generation
-To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance.
-We use the WVS validation set—separated from the training data, which served as the retrieval corpus (as outlined in Section 3)—to evaluate the models.
-Table 4 presents a comparison of our method, using only summaries of values, against four baseline methods.
-Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
-Methods GPT-4o-mini Gemini-1.5-flash
-Zero-Shot 0.6176 0.6041
-Role-Assignment 0.6747 0.6505
-Few-Shot Learning 0.6359 0.6086
-Hybrid Method 0.6670 0.6354
-Values Augmented 0.6894 0.6583
-Table 4: Accuracy comparison between baseline methods and our proposed Values Augmented Generation method using the WVS validation set.
-Bold text indicates the best performance, underlined text indicates the second-best performance.
-This result confirms the effectiveness and robustness of the values-augmented generation approach.
-ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses.
-Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors.
-The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
-6 Conclusion
-We present ValuesRAG, an innovative Retrieval-Augmented Generation framework that revolutionizes cultural values alignment through its comprehensive approach to values context analysis.
-While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles.
-Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines.
-By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses.
-This structured design enables ValuesRAG to bridge the gap between generic LLM capabilities and the demands of culturally sensitive applications, providing a scalable and robust solution for real-world use cases.
-7 Limitation
-Although our baseline comparisons indicate that ValuesRAG generally delivers superior performance compared to alternative methods, it does not always guarantee an exact match to individual’s true values.
-Since we rely on the WVS dataset to summarize individual profiles, there can be mismatches when these summaries are applied to other test sets.
-In future work, we plan to explore more adaptive retrieval strategies that can better align with novel datasets, as well as investigate how integrating additional fine-tuning with retrieval-augmented generation may further refine each agent’s contextual accuracy.
-Appendix A Appendix A. Dataset Description
-A.1 WVS
-As shown in Figure 3, the WVS dataset conducts surveys in over 120 countries worldwide, covering most regions and countries across the globe, making it suitable for constructing a documents corpus for Retrieval Augmented Generation (RAG).
-It is also representative in terms of demographic characteristics.
-We select the most recent WVS survey (Wave 7) to ensure the dataset is as close as possible to the contemporary world.
-Figure 3: Distribution of Country in WVS dataset. The WVS dataset is the most extensive values survey dataset in terms of coverage.
-According to the classification in the codebook released by WVS, the values-related questions it contains can be divided into 13 topics, as shown in Table 5, covering most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values.
-Topic Count
-Social Values, Norms, Stereotypes 45
-Happiness and Wellbeing 11
-Social Capital, Trust and Organizational Membership 47
-Economic Values 6
-Perceptions of Corruption 9
-Perceptions of Migration 10
-Perceptions of Security 21
-Index of Postmaterialism 6
-Perceptions about Science and Technology 6
-Religious Values 12
-Ethical Values 23
-Political Interest and Political Participation 35
-Political Culture and Political Regimes 25
-Table 5: Distribution of Values-related Questions in WVS. The questions were categorized into 13 topics with a total of 259 questions covering most of the dimensions of values
-A.2 Test Datasets
-To validate the robustness of our method, we selected datasets from six different regions as test sets.
-These test sets include both values-related questions and demographic characteristics.
-The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG.
-The values-related questions are utilized as test questions to calculate accuracy (ACC).
-EVS
-The first dataset comes from the European Values Study [^8] https://europeanvaluesstudy.eu, a large-scale, cross-national, and longitudinal survey research program designed to explore values, beliefs, and attitudes across Europe.
-This dataset includes a total of 211 values-related questions and captures 34 demographic characteristics of the respondents.
-We select EVS 2017, conducted in 2017, ensuring alignment with the World Values Survey (WVS) in terms of the timeframe.
-GSS
-We select GSS to represent the population of the United States.
-The General Social Survey [^9] https://gss.norc.org is a sociological survey that has been conducted since 1972 by the National Opinion Research Center (NORC) at the University of Chicago.
-Its primary purpose is to collect and analyze data on the opinions, behaviors, and demographic characteristics of adults in the United States, thereby monitoring societal change and the growing complexity of American society.
-Its questionnaire covers a comprehensive and wide range of topics, including many values-related questions.
-Specifically, within the GSS, we identify 44 questions as values-related and 33 questions as demographic characteristics.
-CGSS
-The Chinese General Social Survey [^10] http://cgss.ruc.edu.cn, initiated in 2003, is China’s earliest national, comprehensive, and continuous academic survey project.
-Conducted by the National Survey Research Center at Renmin University of China, the CGSS systematically collects data at multiple levels, including society, communities, families, and individuals.
-CGSS only provided the questionnaire and data in Chinese, which we have translated into English to ensure its usability.
-We ultimately compile a total of 58 values-related questions and 13 demographic characteristics.
-ISD
-To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set.
-However, we were unable to obtain data from several government surveys in India, so we used data published by the Pew Research Center instead.
-The Pew Research Center’s India Survey Dataset [^11] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance.
-Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others.
-This dataset covers 33 values-related questions and 23 demographic characteristics.
-LAPOP
-The Latin American Public Opinion Project (LAPOP) [^12] https://www.vanderbilt.edu/lapop is a research institute based at Vanderbilt University in Nashville, Tennessee.
-LAPOP’s most notable survey is the AmericasBarometer, the most extensive survey of democratic public opinion and behavior covering the Americas, including North, Central, South America, and the Caribbean.
-This survey measures democratic values and behaviors through voter surveys, providing valuable insights into public sentiments across the region.
-We select this dataset to represent the population of Latin America.
-There are 48 values-related questions and 12 demographic characteristics.
-Africa
-Afrobarometer [^13] https://www.afrobarometer.org is a pan-African, non-partisan research network established in 1999 that conducts public attitude surveys on democracy, governance, economic conditions, and related issues across Africa.
-We selected data from the 8th round of Afrobarometer (collected in 2022), covering 34 African countries.
-After screening, a total of 144 values-related questions and 14 demographic characteristics are obtained.
-Appendix B Appendix B. Prompts Used
-This section provides the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries.
+For the retrieval task, we employed the E5 (base) model (Wang et al.
+2022), which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity.
+To refine this retrieval, we apply a reranker using the GTE-multilingual-reranker-base model (Zhang et al.
+2024).
+Each retrieved candidate is paired with the test individual’s demographic summary and passed as a sentence pair to the reranker.
+The model outputs a relevance score for each pair, which we use to rank all candidates.
+The top-k summaries with the highest scores are then selected as the final set used for generation.
+This two-stage process: retrieval followed by reranking, ensures both semantic similarity and contextual relevance.
+Prompts Used
+We provide the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries below:
 Prompt for Question Answering Task: Respond to the question as the target individual, selecting the answer that aligns with their values and demographic context.
@@
 Based on the provided context, summarize this person’s demographics in one paragraph.
-Appendix C Appendix C. Case Study
-Example 1 Question Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all?
-Original Demographics Summary The respondent is a 63-year-old male born in the United States, of White descent, who speaks English at home.
-He completed 12th grade and holds a high school diploma, while his spouse has the same educational background.
-They are currently married and live in a household of two, with no children.
-The respondent identifies as middle class and reported a family income of \$90,000 to \$109,999 last year, with his own earnings falling within the same range.
-He practices Protestantism and is affiliated with the United Presbyterian Church in the U.S.A.
-His mother has completed two years of college, while his father finished high school.
-The respondent’s demographic background reflects a stable family life with no financial challenges reported.
-Retrieved Summaries: Retrieved Text 1
-Demographics
-The respondent is a 60-year-old White, non-Hispanic male born in the United States, where he speaks English at home.
-He was raised in a family where both parents were also born in the U.S., and he holds a bachelor’s degree, signaling a level of educational attainment typical of the upper middle class.
-He is currently employed full-time in a professional and technical role within the private business sector and is the chief wage earner in his household, which consists of one person.
-The respondent is divorced and has two children.
-He identifies as a Protestant and has reported saving money over the past year, rating his household income as relatively high on a scale of one to ten.
-Values
-This person’s values center on family, personal empowerment, and social equity, emphasizing a balanced life encompassing work, political engagement, and community well-being.
-They prioritize financial responsibility, health, and individual liberties while advocating equality and rejecting traditional gender roles.
-They approach trust cautiously, emphasize accountability in governance, and hold nuanced views on immigration and moral integrity.
-They believe in science and technology’s benefits and maintain personal faith amid diverse beliefs.
-Politically active, they support democratic principles and civil rights while voicing dissatisfaction with current institutions.
-Overall, they blend progressive values focused on social justice, environmental sustainability, and the power of citizen engagement in politics and public life.
-Retrieved Text 2
-Demographics
-The respondent is a 63-year-old White, non-Hispanic male born in the United States, where he speaks English at home.
-He is a citizen of the U.S., with both parents also being born in the country.
-He is married and lives in a household of two people, including himself and his spouse, who is retired.
-He has two children and both he and his spouse completed 12th grade, earning a high school diploma.
-He works full-time in a semi-skilled occupation within the private sector and identifies as belonging to the middle class, reporting that his household income places them in the third income tier.
-The respondent aligns with the Protestant religion and describes their financial situation as stable.
-Values
-This individual values family and work as life’s center, while moderately engaging in politics.
-They emphasize independence, hard work, tolerance, and determination in children, embrace progressive stances on gender and social issues, and place little importance on religious conformity.
-Their trust is selective, relying heavily on family, showing moderate faith in others, and opposing corruption, while valuing economic equality and environmental concerns.
-They uphold science and technology, favor personal freedom over security, and maintain moral integrity and accountability.
-Strongly engaged in civic matters, they are committed to democracy, civil rights, and active political participation, yet remain cautious about government roles.
-Overall, their values balance familial dedication, personal agency, ethical living, and a progressive yet pragmatic outlook on social and political issues.
-Retrieved Text 3
-Demographics
-The respondent is a 65-year-old white male, born in the United States, …
-Example 2 Question How many children to have is a personal matter, and the government should not interfere.
-Do you agree?
-Original Demographics Summary The respondent is a 27-year-old male from China of Han ethnicity, who speaks Chinese at home.
-He holds a postgraduate degree and has never engaged in any work for economic income.
-Currently, he lives alone without any additional household members and identifies as belonging to the middle class, describing his household’s economic status as average in his local area.
-He does not practice any religion and reports no personal income in the past year.
-Retrieved Summaries: Retrieved Text 1
-Demographics
-The respondent is a 28-year-old male of Chinese ethnicity, born in the country where he currently resides.
-He speaks Standard Chinese (Mandarin/Putonghua/Guoyu) at home and identifies as part of the working class.
-He has attained a Bachelor’s degree and is employed full-time in a clerical role within the private sector.
-He lives in a household of four, including his own parents, and is currently married but has no children.
-Both he and his spouse have a similar educational background and are employed full-time.
-His family’s financial situation indicates that they have spent some savings and borrowed money over the past year.
-He does not belong to any religious denomination and has reported his household income as being in the third income level out of ten.
-Values
-This individual holds a complex set of values that place a significant emphasis on family, work, and personal responsibility, while also promoting income equality and valuing hard work as a means of success.
-They have a moderate stance on political engagement, expressing skepticism toward certain institutions while maintaining high trust in organizations like the police and government.
-Their views reflect a critical perspective on issues such as immigration and gender roles, where they value cultural diversity but prioritize native citizens for job opportunities.
-They exhibit a commitment to environmental protection, scientific advancements, and ethical living, while acknowledging their non-religious identity.
-Overall, their values convey a blend of personal autonomy, accountability, and a pragmatic view on governance, demonstrating a preference for democratic principles intertwined with the belief in the importance of social order and national security.
-Retrieved Text 2
-Demographics
-The respondent is a 29-year-old male of Chinese descent, and speaks other Chinese dialects at home.
-He has attained a bachelor’s degree and works full-time in the sales sector within a private business or industry.
-Living alone, he is single and has no children.
-His family background includes parents with upper secondary education.
-Socially, he identifies with the working class and describes his household’s financial status as just getting by, placing his household income in the fourth income group on a scale of ten.
-He does not identify with any religious denomination.
-Values
-This person exhibits a set of values that emphasize the importance of family, hard work, and personal responsibility, reflecting a traditional view of societal roles and a commitment to instilling these qualities in children.
-They hold a pragmatic and somewhat secular worldview, expressing low interest in politics and religion while maintaining a positive attitude towards health, personal agency, and economic stability.
-They advocate for income equality and environmental sustainability, demonstrating concern for social equity and opposing corruption, while also valuing trust primarily within familial ties.
-Additionally, their perspective on immigrants is cautiously appreciative, and they prioritize freedom and economic progress, alongside a recognition of the complexities of modern ethical and moral issues.
-Overall, this individual’s values highlight a balance between familial loyalty, economic and environmental responsibilities, personal freedom, and a discerning approach to broader societal structures.
-Retrieved Text 3
-Demographics
-The respondent is a 25-year-old male of Chinese ethnicity, who speaks Standard Chinese languages, …
+Baseline Methods and Implementation.
+Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach (Tao et al.
+2024), (3) a few-shot learning method (Choenni and Shutova 2024), and (4) a hybrid method that combines both (1) and (2).
+For the role-assignment baseline, we specifically use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data.
+For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts.
+The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set.
+Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance.
+[^10] A detailed analysis of varying $k$ and its implications on model performance is provided in Ablation Study - Varying the Number of Retrieved Summaries.
+Evaluation Metrics.
+We utilize accuracy as the primary evaluation metric, following previous work (Choenni and Shutova 2024), by converting multiple-choice responses into a binary format for consistency and simplicity.
+Specifically, we transform each original response $r_{i}$, ranging on a scale (e.g., Likert scale from 1 to 10), into a binary value $b_{i}$ indicating disagreement or agreement as follows.
+Here, $m$ denotes the midpoint of the response scale:
+$b_{i}=\begin{cases}0,&\text{if }r_{i}\leq m\quad\text{(disagree)}\\[6.0pt] 1,&\text{if }r_{i}>m\quad\text{(agree)}\end{cases}$ (9)
+This categorization metrics effectively captures distinct answer patterns and aligns naturally with values-related questions often posed through Likert-scale formats.
+Model Methods EVS GSS CGSS ISD LAPOP Africa Avg. Accuracy
+GPT-4o mini Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
+Role-Assignment (2024) 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
+Few-Shot Learning (2024) 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
+Hybrid Method 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
+ValuesRAG 0.6021* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
+Gemini 1.5 Flash Zero-shot Inference 0.5419 0.6408 0.4502 0.6017 0.4149 0.4181 0.5113
+Role-Assignment  (2024) 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
+Few-Shot Learning  (2024) 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
+Hybrid Method 0.5845 0.7193 0.5026 0.6253 0.4448 0.5166 0.5655
+ValuesRAG 0.5869 0.7686* 0.5337* 0.6789* 0.4705 0.5473* 0.5977*
+Table 3: Accuracy scores for various methods compared with multiple baselines across six regional datasets.
+$k$ indicates the number of summaries to be retrieved.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
+* denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p<0.05$) over all baseline model(s).
+denotes our proposed method.
+Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg. Accuracy
+GPT-4o mini 1 0.5960 0.7722 0.5347 0.6853 0.4682 0.5905 0.6078
+3 0.6021 0.7781 0.5387 0.7001 0.5030 0.5953 0.6195
+5 0.6052 0.7706 0.5301 0.7016 0.5061 0.5905 0.6174
+10 0.6020 0.7380 0.5317 0.7014 0.5030 0.5680 0.6074
+Gemini 1.5 Flash 1 0.5753 0.7668 0.5272 0.6646 0.4548 0.5369 0.5876
+3 0.5869 0.7686 0.5337 0.6789 0.4705 0.5473 0.5977
+5 0.5868 0.7690 0.5303 0.6734 0.4661 0.5498 0.5959
+10 0.5852 0.7665 0.5279 0.6773 0.4509 0.5464 0.5924
+Table 4: Accuracy scores across six regional datasets for the ablation study: Varying the Number of Retrieved Summaries.
+Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
+Experimental Analysis
+In our experiments, we compare ValuesRAG with four baseline methods: (1) zero-shot inference, (2) role-assignment, (3) few-shot learning, and (4) a hybrid approach that combines role-assignment and few-shot learning.
+Following previous work, as detailed in the Evaluation Metrics section, we specifically evaluate model performance using accuracy.
+All responses are binarized based on contrasting answer patterns to ensure consistency across different values-related questions.
+Experimental results are shown in Table 3.
+We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches.
+By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance.
+Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically.
+Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings.
+As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios.
+The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity.
+However, it still remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
+In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific values-related cultural data for each agent.
+By focusing on values as the primary retrieval targets, this retrieval-augmented framework enables the model to include an expansive set of contextual clues, helping it reflect the depth and breadth of each individual’s background and values.
+Crucially, our ValuesRAG provides a more adaptive mechanism for representing the subtle interplay of personal beliefs and cultural norms via avoiding the limitations of rigid demographic labels or small-sample prompts.
+ValuesRAG more effectively captures the complex dynamics that can shape a respondent’s stance on different questions.
+Evaluations across diverse test datasets demonstrate that ValuesRAG with $k=3$ consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
+Ablation Study
+We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG.
+First, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance.
+We subsequently isolate the effect of using only values-based generation.
+Varying the Number of Retrieved Summaries
+To quantify how the number of retrieved summaries $k$ impacts ValuesRAG’s performance, we evaluate $k\in\{1,3,5,10\}$.
+Table 4 reports accuracy on six regional datasets for both GPT-4o-mini and Gemini 1.5 Flash.
+For GPT-4o-mini, increasing $k$ from 1 to 3 boosts accuracy on five of six datasets, rising from an average of 0.6078 to 0.6195, and yields the best scores on GSS (0.7781), CGSS (0.5387), and Africa (0.5953).
+Although $k=5$ further improves EVS (0.6052) and ISD (0.7016), it degrades GSS and CGSS, dropping the average to 0.6174.
+At $k=10$, performance falls across most datasets (average 0.6074), indicating that an excessive number of summaries dilute relevance.
+Similarly, for Gemini 1.5 Flash, $k=3$ attains the highest overall accuracy (0.5977), with top scores on EVS (0.5869), CGSS (0.5337), ISD (0.6789), and LAPOP (0.4705).
+While $k=5$ slightly edges out $k=3$ on GSS (0.7690) and Africa (0.5498), it lowers EVS and ISD and yields a lower average (0.5959).
+Retrieval at $k=10$ further declines.
+These results reveal a clear trade-off: $k=1$ constrains diversity, whereas $k>3$ introduces marginally relevant or noisy summaries and increases latency.
+Consequently, $k=3$ achieves the best balance between contextual breadth, accuracy gains, and computational cost.
+We thus adopt $k=3$ as our default retrieval depth across all experiments.
+Impact of Values-Only Generation
+To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance.
+We use the WVS validation set: separated from the training data, which served as the retrieval corpus (as outlined in Section: Datasets), to evaluate the models.
+Table 5 presents a comparison of our method, using only summaries of values, against four baseline methods.
+Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
+Methods GPT-4o-mini Gemini-1.5-flash
+Zero-Shot 0.6176 0.6041
+Role-Assignment 0.6747 0.6505
+Few-Shot Learning 0.6359 0.6086
+Hybrid Method 0.6670 0.6354
+Values Augmented 0.6894 0.6583
+Table 5: Accuracy comparison between baseline methods and Values Augmented Generation method using the WVS validation set.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
+Compared to zero‐shot inference (0.6176 for GPT-4o-mini; 0.6041 for Gemini), Values-Only achieves 0.6894 with GPT-4o-mini and 0.6583 with Gemini, demonstrating structured value context provides substantially richer guidance than unconstrained prompts.
+Against the role-assignment (0.6747 for GPT-4o-mini; 0.6505 for Gemini), Values-Only attains higher scores as well, confirming that our generated value summaries capture more nuanced cultural patterns than static demographic labels.
+Similarly, Values-Only outperforms few-shot learning (0.6359 for GPT-4o-mini; 0.6086 for Gemini) and the hybrid method (0.6670 for GPT-4o-mini; 0.6354 for Gemini), further illustrating even without explicit examples or demographic anchors, our values-driven prompts produce more context-aligned predictions.
+These results confirm the effectiveness and robustness of the values-augmented generation approach.
+ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses.
+Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors.
+The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
+Case Study
+To further illustrate the effectiveness and practical utility of ValuesRAG, we present two qualitative case studies derived from the GSS (United States) and CGSS (China) datasets.
+These case studies clearly demonstrate how our method dynamically forms a culturally and demographically relevant retrieval corpus, enhances values alignment, and consistently outperforms baseline approaches by providing richer, context-sensitive reasoning for the generation process.
+Figure 4: Case Study (United States): Cultural Attitudes Toward Premarital Sex.
+The original demographic profile (a 63-year-old Protestant male from the U.S.) is complemented by multiple retrieved summaries, enabling the LLM to reason with contextual sensitivity, avoiding stereotypes and enhancing values alignment.
+Case 1: Cultural Attitudes Toward Premarital Sex (United States)
+The first case (Figure 4) involves a 63-year-old middle-class Protestant male from the U.S., responding to the question: “Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all?” ValuesRAG constructs the retrieval corpus by embedding demographic information and retrieving topically relevant profiles with similar backgrounds (e.g., age, ethnicity, education, and religious affiliation).
+Specifically, the top three retrieved profiles closely align demographically with the original individual yet introduce nuanced ideological variations.
+Each retrieved document not only reflects the target individual’s demographic features but also articulates distinctive value orientations, highlighting the contextual breadth and depth of the retrieved corpus.
+This carefully constructed retrieval corpus significantly enhances values alignment by providing multiple representative and nuanced perspectives, allowing the LLM to consider various contextually valid stances rather than reinforcing stereotypes typically found in static demographic-based methods.
+Compared to role-assignment or few-shot approaches, which struggle to represent this nuanced cultural diversity, ValuesRAG dynamically integrates subtle but critical distinctions across retrieved summaries.
+As a result, the generated response reflects realistic and culturally plausible reasoning that respects both traditional values and evolving societal norms, showcasing the superiority of ValuesRAG in handling culturally sensitive and complex value judgments.
+Figure 5: Case Study (China): Family Planning Preferences.
+The retrieved summaries facilitate nuanced reasoning around personal autonomy and governmental roles, effectively capturing contemporary social shifts in attitudes toward family planning.
+Case 2: Family Planning Preferences (China)
+The second case (Figure 5) examines a 27-year-old secular male with postgraduate education from urban China addressing the question: “How many children to have is a personal matter, and the government should not interfere.
+Do you agree?” In this scenario, ValuesRAG retrieves profiles of individuals with closely matching demographics—such as age, education, urban residence, etc.
+Each retrieved summary reflects values related to personal agency, social responsibility, and family planning, with slight variations in economic status or professional roles.
+The constructed retrieval corpus is well-aligned with the original profile and relevant to the thematic context of the prompt.
+By incorporating these diverse yet demographically coherent value profiles, ValuesRAG enables the LLM to generate responses that account for a variety of socially grounded perspectives on family decision-making.
+In contrast to static prompting methods that may overlook such intra-cultural nuances, ValuesRAG captures shifts in value expression across population segments in a context-aware manner.
+This case shows how the method’s dynamic retrieval mechanism enhances cultural alignment by offering broader coverage of relevant value orientations within the target demographic.
+Conclusion
+We propose ValuesRAG, a novel framework designed to advance cultural values alignment through context-aware reasoning.
+In contrast to prior methods that rely on fixed demographic labels or limited few-shot prompting, ValuesRAG dynamically retrieves and integrates contextually rich value summaries using adaptive retrieval, reranking, and In-Context Learning (ICL).
+Our contributions extend beyond the framework design to include a comprehensive evaluation across diverse, culturally and geographically representative test datasets.
+Extensive experiments demonstrate that ValuesRAG consistently outperforms existing approaches, effectively capturing complex cultural nuances, reducing biases, and generating contextually aligned responses.
+Through its structured and adaptive design, ValuesRAG bridges the gap between general-purpose LLMs and the demands of culturally sensitive applications, offering a scalable and robust solution for real-world deployment.
+Limitation
+Although our baseline comparisons indicate that ValuesRAG generally delivers superior performance compared to alternative methods, it does not always guarantee an exact match to an individual’s true values.
+Specifically, since our method relies on dynamically retrieved summaries from a predefined corpus (WVS), mismatches may occur when summaries derived from one dataset are applied to different or novel datasets.
+Future research should therefore explore more adaptive retrieval strategies capable of precisely aligning with novel datasets and investigate how integrating additional fine-tuning steps within retrieval-augmented generation frameworks could further enhance each agent’s contextual accuracy and generalizability across diverse value domains.
+Ethical Considerations
+While ValuesRAG effectively mitigates the common stereotype reinforcement seen in static methods, its reliance on demographic features inherently carries potential ethical risks related to profiling, bias, and fairness.
+If deployed in sensitive or high-stakes contexts such as public policy or persuasive communication, improper handling of demographic information could unintentionally perpetuate stereotypes or biases, adversely affecting vulnerable groups.
+We emphasize that our framework is intended not to reinforce specific demographic-driven stereotypes or biases, but rather to surface contextually relevant values explicitly for scrutiny.
+Researchers and practitioners utilizing ValuesRAG must proactively examine and address ethical implications surrounding demographic-based retrieval.
+It is critical to integrate ValuesRAG into broader evaluative frameworks designed for continuous monitoring of potential societal impacts, ensuring its responsible deployment.
+Additionally, future work should explicitly investigate the ethical dimensions of demographic profiling within retrieval-augmented methods, further promoting fairness, accountability, and transparency in AI-driven cultural alignment tasks.